OpenStudy (jpes2193):

Challenge time: I have a fractional exponent problem (posted in the comments). Could I get a step by step?

OpenStudy (jpes2193):

\[\frac{ X ^{3/4}Y }{ XY ^{-1/2} }\]

OpenStudy (jhannybean):

\[\frac{x^{3/4}y}{xy^{-1/2}}\] OK... put the \(y^{-1/2}\) in the numerator.

OpenStudy (jhannybean):

\(x^{-1} \iff \dfrac{1}{x}\) Therefore \(\dfrac{1}{x^{-1}} \iff \dfrac{x^1}{1}\)

OpenStudy (jhannybean):

This becomes \(\dfrac{x^{3/4}}{x^1} \cdot (y^1 \cdot y^{1/2})\)

OpenStudy (jhannybean):

Remember, \(x^m \cdot x^n = x^{m+n}\) where m and n are exponents, and x is the base function, or variable.

OpenStudy (jhannybean):

Another rule: \(\dfrac{x^{m}}{x^n} = x^{m-n}\)

OpenStudy (jpes2193):

ok

OpenStudy (jhannybean):

Try working it out with the given formulas, and i'll check your work :)

OpenStudy (jpes2193):

\[\frac{ y ^{3/4} }{ x ^{1/4} }\]

OpenStudy (jhannybean):

almost.

OpenStudy (jpes2193):

?

OpenStudy (jhannybean):

\[y^1 \cdot y^{1/2} = y^{2/2 + 1/2} = y^?\]

OpenStudy (jpes2193):

i wrote 3/2... and typed 3/4...

OpenStudy (jhannybean):

:)

OpenStudy (jpes2193):

its seriously been a day with these problems haha...

OpenStudy (jpes2193):

thank you for all your help :)

OpenStudy (jhannybean):

It's paid off though! You've got this :)

OpenStudy (jpes2193):

i sure hope so the test is tomorrow :) lol

OpenStudy (jhannybean):

Good luck! Do well :)